OpenStudy (anonymous):

the store is having a sale of 20% off a single item. you also have a coupon for $5 off a item. a. write a function to model the cost of an item with 20% discount b. write a model to cost of an item with $5 coupon

OpenStudy (anonymous):

a) y=x-0.2x b) y= x-0.2x -5

OpenStudy (anonymous):

thank you